# Routing algorithm within the multiple non-overlapping paths approach for quantum key distribution networks

### v1

@@ I Introduction
Quantum key distribution (QKD) technologies represent an elegant mechanism for solving the key distribution problem [1, 2, 3, 4], which is one of the central tasks of cryptography. Existing QKD protocols rely on the use of quantum channels for transmitting information that is encoded in quantum states of single photons (on practice, the transmission of weak laser pulses via fiber-optic and free-space communication channels is used) and authentic classical channels [1, 5]. The idea behind is that any interception in the process of transmitting quantum information can be detected by monitoring the quantum bit error rate (QBER) in the quantum channel, while authentication of classical communications is needed for the protection from man-in-the-middle attack during post-processing procedures [1, 2, 6]. In the view of the fact that the security of QKD is based on the laws of quantum physics, it is guaranteed to be secure against any unforeseen technological developments, such as cryptoanalysis using quantum computing [7]. QKD technologies have attracted a great deal of interest and it is a widely studied field of quantum technologies [3, 8]. At the same time, existing realizations of QKD systems still face a number of challenges [8], such as limitations in the key generation rate and distance, and also practical security.
@@
The most developed concept for large-scale, industrial QKD networks is to use trusted nodes [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which play a role of a user in QKD protocol that compose a common secret key by using the bit-wise XOR operation for two established keys with neighboring nodes. The negative side for the trusted-node paradigm is that all the users of the network must trust the node and the node needs to be physically isolated well, while the positive side is reducing the cost and complexity as compared to the all connected point-to-point links. At this stage, the deployment of QKD networks faces the problem of the optimal routing, which has a goal to optimally use the resources of the network when establishing keys between two arbitrary users. Existing QKD networks solve the routing problem by means of various protocols [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], and tools (such as optical switches [38, 21, 26]); however, this problem is far from being fully solved especially in the case of QKD networks of arbitrary topology [31, 32, 33, 34, 35, 36, 37].
In this work, we consider the key routing task within the multiple non-overlapping paths approach [39, 40, 41, 42, 43]. The idea of this approach is to minimize the requirement for trust by distributing a key across remote nodes via independent routes. This ensures that the final key remains secure, provided that all trusted nodes on at least one route remain secure. Importantly, key routing algorithms designed for the standard single-path trusted node schemes do not guarantee the efficiency of key distribution in the context of the non-overlapping paths approach. To address this challenge, we propose and test the first, to our knowledge, key routing algorithm specifically tailored for the non-overlapping paths framework.
@@ II The concept of the $M$–paths scheme
To overcome the limitation of the QKD range, multiple short point-to-point QKD links are combined into QKD networks with so-called intermediate trusted nodes [9, 44, 26]. In this approach, quantum keys are generated only between the network nodes connected directly by a quantum channel, and trusted nodes perform a key hopping via subsequent bit-wise XOR operations. With the use of this technology, it is possible to build QKD networks of various topologies like backbone, ring, star or mixed one.
The downside of the trusted node paradigm is that remote network nodes have to trust the physical integrity of all nodes in the chain between them. It is possible to relax the trust requirement and increase the security by considering an $M$ non-overlapping paths approach [39, 40, 41, 42, 43], where the secret key between two remote nodes is obtained by combining keys from $M$ (with $M>1$) non-overlapping paths. In the following, we refer to it as $M$–paths approach.
@@
where $\bigoplus$ stands for a bit-wise exclusive or (XOR) operation. Therefore, due to the perfect secrecy of the OTP, the eavesdropping becomes technically more complicated in the $M$–path scheme since the intermediate nodes from each of $M$ paths must be hacked (compromised) in order to reconstruct the entire secret key. Equivalently, compromising fewer than $M$ nodes from ${\cal P}_{ij}^{M}$ would not compromise ${\cal K}_{ij}$.
Here, we would like to emphasize several points related to the $M$–paths scheme. First, the applicability of $M$–path schemes strictly depends on the topology of QKD networks. It is impossible to use this scheme in backbone-type networks, where there is only a single path between two remote nodes, or to apply a 3–path scheme within a ring topology of the network graph. Second, distributing an $\ell$–bit key using the $M$–path scheme reduces the remaining keys by $\ell$ bits for each direct link in each path (we discuss this in more detail in the following section). Therefore, increasing $M$ also increases the “cost” of distributing keys between remote nodes in terms of the remaining key between network nodes. Thirdly, the secret key ${\cal K}_{ij}$ can be obtained by concatenating several keys corresponding to different $M$–element sets of non-overlapping paths $\{{\cal P}_{ij}^{M(k)}\}_{k}$:
@@ III Routing algorithm for the $M$–path scheme
The proposed routing Algorithm 1 follows an iterative process. As the basic target quantity to be minimized in each iteration we consider
@@ IV Performance demonstration of the algorithm
To demonstrate the performance of the developed algorithm, we consider the task of routing the key within the $2$–path scheme for two examples of QKD networks. The first QKD network is represented by a six-vertex graph shown in Fig. 2(b). For simplicity, we assume that all adjacent links are equidistantly located and have equal initial key generation rates, namely, $R_{ij}\coloneqq 1$ kbit/s for all $(i,j)\in E$. The target rates are also set to be equal, $T_{ij}\coloneqq 0.1$ kbit/s, for all possible pairs. The decrease of the cost function $\Delta$ with iteration number $r$ is shown in Fig. 2(b). The algorithm completes the task after 80, 160 and 800 iterations, respectively. For all considered values of $\delta R$, the matrix of final effective rates $R^{\text{eff}}$ approaches to a form shown in Fig. 2(c). The corresponding output routing list for remote nodes is obtained in the following:
@@
The second considered example of a more complex QKD network topology, which is taken from Ref. [34], is shown in Fig. 3(b). In contrast to the previous case, the network has different rates for each link and does not possess a symmetrical structure. The secret key generation rates are simulated for the decoy-state BB84 protocol [45, 46, 47] using the link distances from the corresponding graph in Ref. [34]. For this example, we set $T_{ij}\coloneqq 1.0$ kbit/s for all possible pairs of nodes and $M=2$. The behavior of the cost function $\Delta$ with iteration number $r$ for $\delta R\in\{0.1,0.05,0.01\}$ is shown in Fig. 3(b). The algorithm stops after 78, 158, and 793 iterations, respectively. One can notice that $\Delta$ does not approach to zero like in the first example and stops around 0.75 since the chosen target matrix turns out to be overrated for this scheme. We also note the choice of $\delta R$ slightly affects the resulting value of the cost function.
Figure 3: (a) The network graph from Ref. [34] with edges that are labeled with the secret key generation rates in kbps, simulated for the decoy-state BB84 protocol [45, 46, 47]. (b) The iteration number dependence of the cost function. (c) The resulting effective secret key generation rates (obtained with $\delta R=0.01$) for various pairs of network nodes.

### v2

@@ I Introduction
Quantum key distribution (QKD) technologies represent an elegant mechanism for solving the key distribution problem [1, 2, 3], which is one of the central tasks of cryptography. Existing QKD protocols rely on the use of quantum channels for transmitting information that is encoded in quantum states of single photons (on practice, the transmission of weak laser pulses via fiber-optic and free-space communication channels is used) and authentic classical channels [1]. The idea behind is that any interception in the process of transmitting quantum information can be detected by monitoring the quantum bit error rate (QBER) in the quantum channel, while authentication of classical communications is needed for the protection from man-in-the-middle attack during post-processing procedures [1, 2]. In the view of the fact that the security of QKD is based on the laws of quantum physics, it is guaranteed to be secure against any unforeseen technological developments, such as cryptoanalysis using quantum computing [4]. QKD technologies have attracted a great deal of interest and it is a widely studied field of quantum technologies [3, 5]. At the same time, existing realizations of QKD systems still face a number of challenges [5], such as limitations in the key generation rate and distance, and also practical security.
@@
The most developed concept for large-scale, industrial QKD networks is to use trusted nodes [14, 15, 16, 17, 18, 19, 20], which play a role of a user in QKD protocol that compose a common secret key by using the bit-wise XOR operation for two established keys with neighboring nodes. The negative side for the trusted-node paradigm is that all the users of the network must trust the node and the node needs to be physically isolated well, while the positive side is reducing the cost and complexity as compared to the all connected point-to-point links. At this stage, the deployment of QKD networks faces the problem of the optimal routing, which has a goal to optimally use the resources of the network when establishing keys between two arbitrary users. Existing QKD networks solve the routing problem by means of various protocols [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and tools (such as optical switches [33, 18, 34]); however, this problem is far from being fully solved especially in the case of QKD networks of arbitrary topology [26, 27, 28, 29, 30, 31, 32].
In this work, we consider the key routing task within the multiple non-overlapping paths approach [35, 36, 37, 38, 39]. The idea of this approach is to minimize the requirement for trust by distributing a key across remote nodes via independent routes. This ensures that the final key remains secure, provided that all trusted nodes on at least one route remain secure. We propose a routing algorithm that is specifically designed for use with the multiple non-overlapping paths approach. This algorithm differs from previously considered algorithms for single-path routing [21, 22, 24, 29, 31] in that it is tailored to handle the unique requirements of the multiple non-overlapping path scenario. The developed algorithm differs from that proposed in [25] by incorporating the limited capacity for key generation in networks and addressing the routing issue for all nodes simultaneously. At the same time, it differs from the more recent scheme for hybrid-trusted networks [30] in that all nodes in the network are assumed to have equal trust in it. Importantly, key routing algorithms designed for the standard single-path trusted node schemes do not guarantee the efficiency of key distribution in the context of the non-overlapping paths approach. To address this challenge, we propose and test the first, to our knowledge, key routing algorithm specifically tailored for the non-overlapping paths framework.
@@ II The concept of the $M$–paths scheme
To overcome the limitation of the QKD range, multiple short point-to-point QKD links are combined into QKD networks with so-called intermediate trusted nodes [6, 40]. In this approach, quantum keys are generated only between the network nodes connected directly by a quantum channel, and trusted nodes perform a key hopping via subsequent bit-wise XOR operations. With the use of this technology, it is possible to build QKD networks of various topologies like backbone, ring, star or mixed one.
The downside of the trusted node paradigm is that remote network nodes have to trust the physical integrity of all nodes in the chain between them. It is possible to relax the trust requirement and increase the security by considering an $M$ non-overlapping paths approach [35, 36, 37, 38, 39], where the secret key between two remote nodes is obtained by combining keys from $M$ (with $M>1$) non-overlapping paths. In the following, we refer to it as $M$–paths approach. We would like to note that, in this context, we refer to a path as an independent key hopping route. From a physical perspective, such independent paths can be realized even within a single transmission line, where certain nodes may be skipped, as is the case of quantum control-based key distribution networks [41, 42].
@@
where $\bigoplus$ stands for a bit-wise exclusive or (XOR) operation. Therefore, due to the perfect secrecy of the OTP, the eavesdropping becomes technically more complicated in the $M$–path scheme since the intermediate nodes from each of $M$ paths must be hacked (compromised) in order to reconstruct the entire secret key. Equivalently, compromising fewer than $M$ nodes from ${\cal P}_{ij}^{M}$ would not compromise ${\cal K}_{ij}$. Assuming a probability of compromise of a node operated as a trusted node is $\epsilon_{\rm compr}$, we can estimate that the probability of compromising a key between any two remote nodes in $M$–paths regime is upper bounded by $\epsilon^{M}_{\rm compr}$.
Here, we would like to emphasize several points related to the $M$–paths scheme. First, the applicability of $M$–path schemes strictly depends on the topology of QKD networks. It is impossible to use this scheme in backbone-type networks, where there is only a single path between two remote nodes, or to apply a 3–path scheme within a ring topology of the network graph. Second, distributing an $\ell$–bit key using the $M$–path scheme reduces the remaining keys by $\ell$ bits for each direct link in each path (we discuss this in more detail in the following section). Therefore, increasing $M$ also increases the “cost” of distributing keys between remote nodes in terms of the remaining key between network nodes. In this regard, the selection of $M$ within the range allowed by the topology should consider a balance between the desired security level, estimated as $\epsilon_{\rm compr}^{M}$, and the increased key usage in the network. Thirdly, the secret key ${\cal K}_{ij}$ can be obtained by concatenating several keys corresponding to different $M$–element sets of non-overlapping paths $\{{\cal P}_{ij}^{M(k)}\}_{k}$:
@@ III Routing algorithm for the $M$–path scheme
Here, we would like to draw attention to the fact that all OTP-secured communications must be authenticated, which also requires the use of a secret key. Due to the fixed consumption of secret keys in authentication of each message, it is more efficient in practice to increase the interval between “key distribution rounds” and, consequently, the length of the keys transmitted in each message. This approach helps to minimize the relative costs associated with authentication. In the following, we assume that the key consumption for authentication can be neglected compared to the transmitted key material.
The proposed routing Algorithm 1 follows an iterative process (see also an example of a step-by-step workflow in Appendix A). As the basic target quantity to be minimized in each iteration we consider
@@
Finally, we would like to discuss the role of two hyperparameters: $r_{\max}$ and $\delta R$. By bounding the number of iterations, $r_{\max}$ limits the maximum running time of the algorithm. One can also consider replacing the condition $r<r_{\max}$ in the main loop of the algorithm by a condition that checks if the current running time is less than the maximum allowed time. If there are no limits on running time, $r_{\max}$ can be set to infinity. $\delta R$ determines the size of the step at each iteration. A smaller $\delta R$ results in more iterations and a higher precision in the final value of the cost function ($\Delta$). As we will see in the testing in the next section, reducing $\delta R$ can improve the cost function by the amount of $\delta R$. However, from a practical perspective, it is not advisable to reduce $\delta R$ below a level corresponding to the typical cost of a secret key for authentication purposes, assumed to be negligible within the construction of the algorithm.
@@ IV Performance demonstration of the algorithm
To demonstrate the performance of the developed algorithm, we consider the task of routing the key within the $2$–path scheme for two examples of QKD networks.
The first QKD network is represented by a six-vertex graph shown in Fig. 2(b). For simplicity, we assume that all adjacent links are equidistantly located and have equal initial key generation rates, namely, $R_{ij}\coloneqq 1$ kbit/s for all $(i,j)\in E$. The target rates are also set to be equal, $T_{ij}\coloneqq 0.1$ kbit/s, for all possible pairs. The decrease of the cost function $\Delta$ with iteration number $r$ is shown in Fig. 2(b). The algorithm completes the task after 80, 160 and 800 iterations, respectively. For all considered values of $\delta R$, the matrix of final effective rates $R^{\text{eff}}$ approaches to a form shown in Fig. 2(c). The corresponding output routing list for remote nodes is obtained in the following:
@@
The second considered example of a more complex QKD network topology, which is taken from Ref. [29], is shown in Fig. 3(b). In contrast to the previous case, the network has different rates for each link and does not possess a symmetrical structure. The secret key generation rates are simulated for the decoy-state BB84 protocol [43, 44, 45, 46, 47] using the link distances from the corresponding graph in Ref. [29]. For this example, we set $T_{ij}\coloneqq 1.0$ kbit/s for all possible pairs of nodes and $M=2$. The behavior of the cost function $\Delta$ with iteration number $r$ for $\delta R\in\{0.1,0.05,0.01\}$ is shown in Fig. 3(b). The algorithm stops after 78, 158, and 793 iterations, respectively. One can notice that $\Delta$ does not approach to zero like in the first example and stops around 0.75 since the chosen target matrix turns out to be overrated for this scheme. We also note the choice of $\delta R$ slightly affects the resulting value of the cost function.
Figure 3: (a) The network graph from Ref. [29] with edges that are labeled with the secret key generation rates in kbps, simulated for the decoy-state BB84 protocol using the simple theoretical model from Ref. [44] and taking into account the detector’s dead time effect. (b) The iteration number dependence of the cost function. (c) The resulting effective secret key generation rates (obtained with $\delta R=0.01$) for various pairs of network nodes.
@@ V Conclusion and outlook
As a potential limitation on the applicability of the developed algorithm, we would like to emphasize its strong dependence on network topology. Specifically, it is assumed that any two remote nodes can be connected by at least $M$ non-overlapping paths. However, in a real QKD network, this assumption may not hold. It is possible to envision a scenario where some pairs of remote nodes are connected by non-overlapping paths, but others are not. This limitation can be addressed by considering a more flexible approach with different requirements for the value of $M$ for each pair, or by considering “partially overlapping path” scenarios. This appears to be a promising avenue for future research.
@@ Acknowledgements.
Appendix A
To illustrate a step-by-step workflow of the Algorithm 1, we consider a simple network of five nodes, depicted in Fig. 4(a). The graph edges are labeled with corresponding secret key generation rates which are represented by the initial key rate matrix in Fig. 4(b). The target rates are assumed to be $T_{ij}=0.2$. The iteration step is set to be $\delta R=0.1$.
Figure 4: (a) The network graph with edges, labeled with secret key generation rates. (b) The initial key rate matrix. (c)–(f) The resulting effective secret key generation rates after every iteration.
  1. 1.
    At the start the remote nodes apparently have the “worst” node pair deficiency,
    $\operatorname*{arg\,max}_{i,j\in V}D(i,j)=\{(0,4),(1,3)\}\,.$
    Hence, a random candidate, e.g. $(1,3)$, is chosen. Then the list of all two non-overlapping path combinations from node 1 to node 3 is constructed,
    ${\cal L}^{2}_{13}=[$$\{(1,0,2,3),(1,4,3)\}\,,$ $\textbf{D}=-0.2$
    $\{(1,0,3),(1,4,2,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,4,3)\}\,,$ $\textbf{D}=-0.2$
    $\{(1,0,3),(1,2,4,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,2,3)\}\,,$ $\textbf{D}=-0.3$
    $\{(1,4,3),(1,2,0,3)\}\,,$ $\textbf{D}=-0.2$
[…]
    Here, in the right column, we put values of ${\bf D}$ for each of combination of paths. It is computed as a maximal defficiecy (target rate minus current rate) over all links in the both paths. For example, for $\{(1,0,3),(1,4,3)\}$, we have ${\bf D}=\max(0.2-0.5,0.2-0.5,0.2-0.4,0.2-0.6)=-0.2$. The lowest value of the path deficiency function, $\textbf{D}=-0.3$, determines the optimal 2–path combination,
    ${\cal P}_{13}^{2,\text{opt}}=\{(1,0,3),(1,2,3)\}\,.$
    The effective rates for all node pairs from ${\cal P}_{13}^{2,\text{opt}}$ are computed as follows:
    $R_{13}^{\text{eff}}$ $=\delta R\,,$
    $R_{ij}^{\text{eff}}$ $=R_{ij}-\delta R\leavevmode\nobreak\ \leavevmode\nobreak\ \forall ij\in\{01,03,12,23\}\,.$
    The result can be seen in Fig. 4(c). The routing list, which was initially empty, is appended with the first entry:
    ${\sf routing\_list}=[\{(1,0,3),(1,2,3)\}:0.1]\,.$
  2. 2.
    After the first iteration, $(0,4)$ automatically becomes the “worst” pair, for which the list of non-overlapping paths with corresponding 2–path deficiencies is
    ${\cal L}^{2}_{04}=[$$\{(0,1,4),(0,2,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,1,4),(0,2,3,4)\}\,,$ $\textbf{D}=-0.2$
    $\{(0,1,4),(0,3,4)\}\,,$ $\textbf{D}=-0.2$
    $\{(0,1,4),(0,3,2,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,1,2,4),(0,3,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,2,4),(0,3,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,2,1,4),(0,3,4)]\,,$ $\textbf{D}=-0.2$
    One can see that there are three candidates with lowest value $\textbf{D}=-0.2$ which form the set ${\cal P}_{04}^{2,\text{pre-opt}}$, but only one has the lowest distance (i.e. the total number of edges that form two paths), equal to 4. Hence,
    ${\cal P}_{04}^{2,\text{opt}}=\{(0,1,4),(0,3,4)\}\,.$
    The effective rates for all node pairs from ${\cal P}_{04}^{2,\text{opt}}$ are recomputed as follows:
    $R_{04}^{\text{eff}}$ $=\delta R\,,$
    $R_{01(03)}^{\text{eff}}$ $\to R_{01(03)}^{\text{eff}}-\delta R\,,$
    $R_{14(34)}^{\text{eff}}$ $=R_{14(34)}-\delta R\,.$
    The result can be seen in Fig. 4(d). The routing list is appended with the second entry:
    ${\sf routing\_list}=[\{(1,0,3),(1,2,3)\}$ $:0.1\,,$
    $\{(0,1,4),(0,3,4)\}$ :0.1]:0.1]\,.: 0.1 ] .
  3. 3.
    After the second iteration there are again two candidates with “worst” node pair deficiency, $\operatorname*{arg\,max}_{i,j\in V}D(i,j)=\{(0,4),(1,3)\}$, among which, e.g. $(1,3)$, is chosen. The non-overlapping path list is apparently the same as for the first iteration, however, the deficiencies are renewed:
    ${\cal L}^{2}_{13}=[$$\{(1,0,2,3),(1,4,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,4,2,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,4,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,2,4,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,0,3),(1,2,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,4,3),(1,2,0,3)\}\,,$ $\textbf{D}=-0.1$
    $\{(1,4,3),(1,2,3)]\,,$ $\textbf{D}=-0.1$
    Among three candidates with lowest distance a random one is chosen:
    ${\cal P}_{13}^{2,\text{opt}}=\{(1,4,3),(1,2,3)\}\,.$
    The new effective rates are
    $R_{13}^{\text{eff}}$ $\to R_{13}^{\text{eff}}+\delta R\,,$
    $R_{ij}^{\text{eff}}$ $\to R_{ij}^{\text{eff}}-\delta R\leavevmode\nobreak\ \leavevmode\nobreak\ \forall ij\in\{14,34,12,23\}\,.$
    The result can be seen in Fig. 4(e). The routing list is appended with new entry:
    ${\sf routing\_list}=[\{(1,0,3),(1,2,3)\}$ $:0.1\,,$
    $\{(1,4,3),(1,2,3)\}$ $:0.1\,,$
    $\{(0,1,4),(0,3,4)\}$ :0.1]:0.1]\,.: 0.1 ] .
  4. 4.
    After the third iteration $D(0,4)=0.1$ turns out to be the highest. Among all paths from node 0 to node 4,
    ${\cal L}^{2}_{04}=[$$\{(0,1,4),(0,2,4)\}\,,$ $\textbf{D}=0$
    $\{(0,1,4),(0,2,3,4)\}\,,$ $\textbf{D}=0$
    $\{(0,1,4),(0,3,4)\}\,,$ $\textbf{D}=0$
    $\{(0,1,4),(0,3,2,4)\}\,,$ $\textbf{D}=0$
    $\{(0,1,2,4),(0,3,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,2,4),(0,3,4)\}\,,$ $\textbf{D}=-0.1$
    $\{(0,2,1,4),(0,3,4)]\,,$ $\textbf{D}=0$
    there are two candidates with lowest 2–path deficiency $\textbf{D}=-0.1$. The optimal path combination with the lowest distance is
    ${\cal P}_{04}^{2,\text{opt}}=\{(0,2,4),(0,3,4)\}\,,$
    In this way the final rate correction is
    $R_{04}^{\text{eff}}$ $\to R_{04}^{\text{eff}}+\delta R\,,$
    $R_{ij}^{\text{eff}}$ $\to R_{ij}^{\text{eff}}-\delta R\leavevmode\nobreak\ \leavevmode\nobreak\ \forall ij\in\{02,24,03,34\}\,.$
    The result can be seen in Fig. 4(e). The resulting routing list takes the form:
    ${\sf routing\_list}=[\{(1,0,3),(1,2,3)\}$ $:0.1\,,$
    $\{(1,4,3),(1,2,3)\}$ $:0.1\,,$
    $\{(0,1,4),(0,3,4)\}$ $:0.1\,,$
    $\{(0,2,4),(0,3,4)\}$ :0.1]:0.1]\,.: 0.1 ] .
    Since now all effective rates are not less than the target rate, the algorithm stops.